(a) An electron has a kinetic energy of 4.10 ev. Find its wavelength. 5905 10 Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully Ans (b) A photon has energy 4.10 ev. Find its 303 103 x is off by a multiple of ten, nm s Previous Arwwers SerCP1027 P033 0075 points In the ground state of hydrogen, the uncertainty in the position of the electron is roughly o 10 nm. irthe speed of the electron isapproximately the same as the uncertainty n its speed, about how fast is dmers from the correct answer by more than 10 Double check your calculations m/s Your response 235 points Pevkus Answers sercP 02TAE 001
Answer #1

Use the DeBroglie relation

\lambda = \frac{h}{p} = \frac{h}{\sqrt{2mK.E}}

a] For the electron: K.E = 4.10 eV

so, \lambda = \frac{6.626\times 10^{-34}}{\sqrt{2(9.1\times 10^{-31})(4.1\times 1.6\times 10^{-19})}} = 6.064\times 10^{-10}m

b] For the proton: K.E = 4.10 eV

so, \lambda = \frac{6.626\times 10^{-34}}{\sqrt{2(1.67\times 10^{-27})(4.1\times 1.6\times 10^{-19})}} = 1.415\times 10^{-11}m

2]

Use Heisenberg's uncertainty relation:

\Delta x \Delta p \geq \hbar

=> \Delta x [m\Delta v ] \geq \hbar

=> v\approx \Delta v = \frac{\hbar}{m\Delta x }

=> v\approx \frac{1.05456\times 10^{-34}}{(9.1\times 10^{-31})0.1\times 10^{-9} }=1.16\times 10^6 m/s or about 106 m/s.